# Evaluation of LLMs in retrieving food and nutritional context for RAG systems

### 10 Mar 2026 v1

@@ I Introduction
To bridge this gap, Retrieval-Augmented Generation (RAG) systems offer a promising alternative by enabling direct natural language querying of databases [3]. In this study, we evaluate how reliably this paradigm operates when applied to food composition data. One critical component of a RAG system is its ability to accurately and reliably retrieve context from a specialized database. Unlike traditional approaches, our system uses a Large Language Model (LLM) to transform a natural language query, such as ”Which foods have more than 12g of protein?” into the structured metadata filters required to efficiently query the underlying Chroma [4] vector database which we use to store and retrieve semantic embeddings, created from Slovenian food composition data and its corresponding metadata. To ensure accessibility for non-technical users, this retrieval mechanism is exposed through a simple user interface that allows food compilers and nutritionists to interact with the system using natural language.
@@ II Related Work
II-A Applications of RAG and LLMs in Food and Nutrition
@@ III-A System Overview
III-B Data Collection and Preprocessing
@@
III-C Metadata Filter Generation
@@ III-D Retrieval Mechanism
III-E Similarity Threshold for Context Selection
@@ III-F Evaluation of the System
III-G Question Difficulty Categorization
@@
III-H Retrieval Mechanism Testing
@@
IV Results
Table I reports the average F1 scores (computed over five independent runs) for all evaluated models across the three difficulty categories and for each of the three similarity thresholds ($\mu-\sigma$, $\mu$, and $\mu+\sigma$).
@@ IV Results
TABLE I: Average F1 scores.

### 11 Mar 2026 v2

@@ 1 Introduction
To bridge this gap, Retrieval-Augmented Generation (RAG) systems offer a promising alternative by enabling direct natural language querying of databases [3]. In this study, we evaluate how reliably this paradigm operates when applied to food composition data. One critical component of a RAG system is its ability to accurately and reliably retrieve context from a specialized database. Unlike traditional approaches, our system uses a Large Language Model (LLM) to transform a natural language query, such as "Which foods have more than 12g of protein?" into the structured metadata filters required to efficiently query the underlying Chroma [5] vector database which we use to store and retrieve semantic embeddings, created from Slovenian food composition data and its corresponding metadata. To ensure accessibility for non-technical users, this retrieval mechanism is exposed through a simple user interface that allows food compilers and nutritionists to interact with the system using natural language.
@@ 2 Related Work
2.1 Applications of RAG and LLMs in Food and Nutrition
@@ 3.1 System Overview
3.2 Data Collection and Preprocessing
@@
3.3 Metadata Filter Generation
@@ 3.4 Retrieval Mechanism
3.5 Similarity Threshold for Context Selection
@@ 3.6 Evaluation of the System
3.7 Question Difficulty Categorization
@@
3.8 Retrieval Mechanism Testing
@@
4 Results
Table 1 reports the average F1 scores (computed over five independent runs) for all evaluated models across the three difficulty categories and for each of the three similarity thresholds ($\mu-\sigma$, $\mu$, and $\mu+\sigma$).
@@ 4 Results
Table 1: Average F1 scores.
@@ 5.1 Limitations
Acknowledgments and Disclosure of Funding
The authors acknowledge the support of the Slovenian Research Agency through program grant No. P2-0098 and project grant No. GC-0001. We also acknowledge the support of the EC/EuroHPC JU and the Slovenian Ministry of HESI via the project SLAIF (grant number 101254461).